OpenStudy (anonymous):

Find an equation of the line containing the given pair of points: (-5, -7) and (-3, -8). The equation of the ine in slope-intercept form is y = ?

OpenStudy (aroub):

slope= -8-7/-3-5 = -15/-8 =15/8 y=mx+b , m=15/8 , (-5,-7) -7=15/8(-5)+b -7=-75/8 +b -56/7+75/8=b 19/8=b so the equation is y=15/8 +19/8